OpenStudy (anonymous):

describe how the graph of y+x^2 can be transformed to the graph of the given equation. y=x^2-12 @ganeshie8 @hero @myininaya @pooja195 @michele_laino @campbell_st @solomonzelman @loser66 @ligi0210 @mathstudent55 @e.mccormick

OpenStudy (anonymous):

can u show me step by step @Loser66

OpenStudy (mathstudent55):

If you have any function, such as y = f(x), and you add a constant to the right side, and the function becomes y = f(x) + k, then the graph shifts k units vertically.

OpenStudy (anonymous):

so it shifts up 12 units right? @mathstudent55

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

can u just tell me if I'm right @Loser66

OpenStudy (mathstudent55):

In your case, k = -12, so it shifts 12 units down.

OpenStudy (anonymous):

@mathstudent55 can you help me with a quick question please :)

OpenStudy (mathstudent55):

y = x^2 compared to y = x^2 - 12 Your function can be written as: y = x^2 + (-12) Since k = -12, a negative number, the shift is down 12 units.

OpenStudy (michele_laino):

hint: we have these graphs: